OpenStudy (anonymous):

find the tangents of the curve 4x^2 + 9y^2 + 24x - 18y -180 = 0 from (-6, 12). the answers are 8x + 9y - 60 = 0 and 16x - 7y + 180. i already have the process but i didnt derived to the answer. please help.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

another messy algebra problem with this one. you should get \[ \frac{dy}{dx}= \frac{-4(x+3)}{9(y-1) } \] and the "slope equation" \[ \frac{-4(x+3)}{9(y-1) } = \frac{(y-12)}{(x+6)}\] after cross multiplying \[ -4(x+3)(x+6) = 9(y-1)(y-12) \] you should be able to manipulate that equation into the form \[ 4x^2 +9y^2= -36x +117y -180 \] to simplify that we use the original equation \[ 4x^2 + 9y^2 + 24x - 18y -180 = 0 \] and re-arrange to get \[4x^2 + 9y^2 = - 24x + 18y +180\]

OpenStudy (phi):

now substitute into the "cross multiply" equation \[4x^2 +9y^2= -36x +117y -180 \\ \text{ to get } \\ −24x+18y+180= -36x +117y -180 \] solve for x or y... I solved for y to get \[ y = \frac{4x+120}{33} \]

OpenStudy (phi):

Use that value of y in the "cross multiply" equation \[ -4(x+3)(x+6) = 9(y-1)(y-12) \\ \text{ to get }\\ -4(x+3)(x+6) = 9\left(\frac{4x+120}{33}-1\right)\left(\frac{4x+120}{33}-12\right) \]

OpenStudy (phi):

let's just concentrate on the right side \[ 9\left(\frac{4x+120}{33}-1\right)\left(\frac{4x+120}{33}-12\right) \\ \cancel{9}3\left(\frac{4x+120-33}{\cancel{33}11}\right)\left(\frac{4x+120-396}{33}\right) \\ \cancel{3}\left(\frac{4x+87}{11}\right)\left(\frac{4x-276}{\cancel{33}11}\right) \\ \frac{4}{121}(4x+87)(x-69) \]

OpenStudy (phi):

the entire equation, using that new right hand side is \[ -4(x+3)(x+6) = \frac{4}{121}(4x+87)(x-69) \] divide by 4 , and multiply by 121 to get \[ -121(x+3)(x+6)= (4x+87)(x-69) \]

OpenStudy (phi):

if you work on that, you should eventually get \[ 5x^2 +36x-153=0 \] one of its roots is x=3 (so (x-3) divides evenly into it) once you have the two values for x (the 2 roots) you can find y and then the slope and finally the equation of the tangent line through point (-6,12)

OpenStudy (anonymous):

gaaah i am so stupid. i messed up in the cross multiplying stage.. thank you so much @phi :) thank you thank you

OpenStudy (phi):

These problems require a LOT of careful work. btw, here is a graph of the problem

OpenStudy (anonymous):

last problem..

OpenStudy (anonymous):

please? very last problem about this topic

OpenStudy (anonymous):

thanks thanks :)